Karen saves her change and has dimes, nickels, and pennies totaling $4.52. If there is 1 more dime than three times the number of nickels and 10 more pennies than nickels, how many of each type of coin are there?
I start the formula like this:
(3x+10)+(x+10)+x=4.52
x is the number of nickels.
3x+10 is 3 times the number of nickels plus one dime.
x+10 is 10 more than the number of pennies.
x is the number of nickels
I've tried to solve this by changing around the decimals n' such but the only time the number comes out even is by dividing both sides by 4 (4x=432--4x/4=432/4) Otherwise they come out with decimals equaling 3-4 places.
Thank you so much :)
Any help is appreciated :)

Answer by mananth(16949)

Karen saves her change and has dimes, nickels, and pennies totaling $4.52. If there is 1 more dime than three times the number of nickels and 10 more pennies than nickels, how many of each type of coin are there?
...
let nickels = x
pennies = x+10
dimes= 3x+1
...
5x+x+10+10(3x+1)=452
5x+x+10+30x+10=452
36x+20=452
36x=452-20
36x=432
/36
x= 432/36
x=12 the number of nickels
pennies = 10+12 = 22
dimes =3x+1 = 37
...
CHECK
12*5 + 22*1+37*10= 452